For any natural number n, prove the validity of given series by mathematical induction:

2(√(n+1)-1)<1+(1/√2)+⋯..+(1/√n)<2√n?


1
Expert's answer
2021-07-19T15:11:58-0400

"2\left( \sqrt{n+1}-1
\right)<1+\tfrac{1}{\sqrt{2}}+…+\tfrac{1}{\sqrt{n}}<2\sqrt{n}"


Base Case: "n=1"

"2\left(\sqrt{2}-1\right)<1<2"

"\sqrt{2}-1<\tfrac{1}{2}<1"

"\sqrt{2}\approx 1.414<1.5<2"


Assume: "2\left( \sqrt{k+1}-1
\right)<1+\tfrac{1}{\sqrt{2}}+…+\tfrac{1}{\sqrt{k}}<2\sqrt{k}"

Prove: "2\left( \sqrt{k+2}-1
\right)<1+\tfrac{1}{\sqrt{2}}+…+\tfrac{1}{\sqrt{k+1}}<2\sqrt{k+1}"


1) "1+\tfrac{1}{\sqrt{2}}+…+\tfrac{1}{\sqrt{k}}+\tfrac{1}{\sqrt{k+1}}<2\sqrt{k}+\tfrac{1}{\sqrt{k+1}}=\frac{2\sqrt{k(k+1)}+1}{\sqrt{k+1}}"


Using inequality "\sqrt{xy}\leq \frac{x+y}{2}" , we get "\frac{2\sqrt{k(k+1)}+1}{\sqrt{k+1}}\leq \frac{(2k+1)+1}{\sqrt{k+1}}=2\sqrt{k+1}"


Therefore, "1+\tfrac{1}{\sqrt{2}}+…+\tfrac{1}{\sqrt{k+1}}<2\sqrt{k+1}"


2) "1+\tfrac{1}{\sqrt{2}}+…+\tfrac{1}{\sqrt{k}}+\tfrac{1}{\sqrt{k+1}}>2\left(\sqrt{k+1}-1\right)+\tfrac{1}{\sqrt{k+1}}=\frac{2(k+1)+1}{\sqrt{k+1}}-2=\frac{2k+3}{\sqrt{k+1}}-2"


Using inequality "\sqrt{xy}\leq \frac{x+y}{2}" , we get "\sqrt{(k+1)(k+2)}\leq \frac{2k+3}{2}" .


So, "2\sqrt{k+2}\leq \frac{2k+3}{\sqrt{k+1}}" .

Therefore, "1+\tfrac{1}{\sqrt{2}}+…+\tfrac{1}{\sqrt{k+1}}>2\sqrt{k+2}-2=2\left(\sqrt{k+2}-1\right)"